OpenStudy (amyna):

Summation from 0 to infinity K+1/3^k Do I use the ratio test? But how do I do it?

OpenStudy (jango_in_dtown):

what is the question?

OpenStudy (jango_in_dtown):

@amyna

OpenStudy (amyna):

It's K+1/3^k

OpenStudy (jango_in_dtown):

convergence/ divergence test?

OpenStudy (jango_in_dtown):

\[k+\frac{ 1 }{ 3^k }\]

OpenStudy (jango_in_dtown):

this one??

OpenStudy (amyna):

It says if this converges absoulty, conditionally, or just diverged

OpenStudy (amyna):

No k+1 is in the numerator

OpenStudy (jango_in_dtown):

\[\sum_{0}^{\infty} \frac{ k+1 }{ 3^k }\]

OpenStudy (amyna):

Yes

OpenStudy (jango_in_dtown):

do the ratio test first...

OpenStudy (jango_in_dtown):

t_k= (k+1)/3^k and t_k+1= (k+2)/3^(k+1)

OpenStudy (amyna):

Oh okay got it! I was just confused if I had to add the k+1 in the numerator or not

OpenStudy (amyna):

Thanks!!

OpenStudy (jango_in_dtown):

replace k by k+1 to get t_k.. so whats the solution? do you need more help?

OpenStudy (jango_in_dtown):

its convergent :)

OpenStudy (amyna):

The numerator becomes k+2 right?

OpenStudy (jango_in_dtown):

which form do you use ? \[\lim |\frac{ a_n }{ a_{n+1} }| \]

OpenStudy (jango_in_dtown):

or \[\lim |\frac{ a _{n+1} }{ a_n }|\] ?

OpenStudy (jango_in_dtown):

@amyna

OpenStudy (amyna):

The second one

OpenStudy (jango_in_dtown):

ok \[\lim |\frac{ t _{k+1} }{ t_k }|=\lim |\frac{ (k+2). 3^k }{ (k+1) .3^{k+1}}|\]

OpenStudy (jango_in_dtown):

= \[\lim |\frac{ 1 }{ 3 }. \frac{ 1+1/k }{ 1+2/k }|=1/3\]

OpenStudy (jango_in_dtown):

<1 so by ratio test it is ?????

OpenStudy (jango_in_dtown):

@amyna

OpenStudy (jango_in_dtown):

in the numerator it will be 1+2/k and in the denominator 1+1/k